OpenStudy (pradius):

Let u and v=24 Check if ||u||+||v|| = ||u+v||

OpenStudy (jdoe0001):

well, I'm assuming v = 24, is written in a+bi form, that is v = 24 + 0i, or <24, 0 >

OpenStudy (pradius):

Im not sure about that... the problem is written just like I did

OpenStudy (jdoe0001):

I'd think is 2 vectors, one written in an ordered pair, and the other as a complex form, either way a + bi = <a, b>

OpenStudy (jdoe0001):

as far as the ||u|| that, is the "magnitude" of the vector is really just \(\huge ||v|| = \sqrt{a^2+b^2}\)

OpenStudy (jdoe0001):

and vector addition is just <a, b> + <c, d> => <a+c, b+d>

OpenStudy (jdoe0001):

so, ||<a, b>|| + ||<24, 0>|| => \(\sqrt{a^2+b^2} + \sqrt{24}\)

OpenStudy (jdoe0001):

||u+v|| that is => || <a,b> + <24, 0> || => || <a+24, b> || => \(\sqrt{(a+24)^2+b^2}\\ \implies \sqrt{a^2+48a+24^2+b^2}\)

OpenStudy (jdoe0001):

so, I gather the answer, they're \(\ne\)

OpenStudy (jdoe0001):

ohh and so, ||<a, b>|| + ||<24, 0>|| => \(\sqrt{a^2+b^2} + \sqrt{24^2}\)

OpenStudy (anonymous):

Let v be <c,d>\(\sqrt{a^2+b^2}+24 \neq \sqrt{(a+c)^2+(b+d)^2}=\sqrt{a^2+2ac+c^2+b^2+2bd+d^2}=\) \(=\sqrt{a^2+2(ac+bd)+b^2 +24^2}\)

OpenStudy (anonymous):

@Pradius

OpenStudy (pradius):

gimme a sec